# The power-spectrum tensor in steady-state systems and its role in quantum friction

### v1

@@ Appendix A Positive definite matrices
In this appendix we elaborate on the reasons for why some of the matrices encountered in the main text are positive semidefinite or positive definite.
Let us consider first the power-spectrum tensor. For a generic complex vector $\mathbf{u}$ and any real-valued function $f(t)$ we can define the operator
[…]
@@
[…]
Then, since $|f(\omega)|^{2}\geq 0$, one must have that
@@
For discussing the properties of $\underline{\alpha}_{\Im}(\omega)$ we consider the perturbed Hamiltonian $\hat{H}^{\prime}_{t}=\hat{H}_{0}-\hat{\mathbf{d}}\cdot\mathbf{e}(t)$, where $\mathbf{e}(t)$ denotes a small ($\left|\mathbf{e}(t)\right|\ll 1$) real three-dimensional vector which is different from zero for a finite interval of time $\mathbf{e}(t\to\pm\infty)\to 0$. In linear order of perturbation theory, the expectation value of the associated time evolution of the dipole operator as prescribed by $\hat{H}^{\prime}_{t}$, i.e. $\hat{\mathbf{d}}^{\prime}(t)$, can be written as
@@
$\langle\hat{\mathbf{d}}^{\prime}(t)\rangle\equiv\mathrm{tr}[\hat{\mathbf{d}}^{\prime}(t)\hat{\rho}]$ $\stackrel{{\scriptstyle\left|\mathbf{e}(t)\right|\ll 1}}{{\approx}}\int_{-\infty}^{\infty}\mathrm{d}t_{1}\;\underline{\alpha}(t-t_{1})\cdot\mathbf{e}(t_{1})~{},$ (51)
@@
where $\underline{\alpha}(\tau)$ is the linear response tensor with elements defined in Eq. (6). Mathematically, we may regard the tensor as the functional derivative with respect to the external perturbation of the exact time evolution, i.e. $\delta\langle\hat{d}^{\prime}_{i}(t)\rangle/\delta e_{j}(t^{\prime})=\alpha_{ij}(t-t^{\prime})$.
@@
In general, the time evolution according to $\hat{H}^{\prime}_{t}$ of an operator $\hat{A}_{t}$ which explicitly depends on time is given by $\hat{A}_{t}(t)=U^{{\dagger}}_{t}\hat{A}_{t}U_{t}$,. The operator $U_{t}$ is the solution of $\partial_{t}U_{t}=(\mathrm{i}/\hbar)\hat{H}_{t}\hat{U}_{t}$ with the initial condition $U_{t=0}=\underline{\mathds{1}}$. Accordingly, the total time derivative of $\hat{A}_{t}(t)$ is given by
$\frac{d}{dt}\hat{A}_{t}(t)=\frac{\mathrm{i}}{\hbar}U^{{\dagger}}_{t}[\hat{H}_{t},\hat{A}_{t}]U_{t}+U^{{\dagger}}_{t}(\partial_{t}\hat{A}_{t})U_{t}~{}.$ (52)
@@
As a consequence, during the perturbation, the total change of energy is given by
[…]
@@
$=-\int_{-\infty}^{\infty}\mathrm{d}t\;\langle\hat{\mathbf{d}}^{\prime}(t)\rangle\cdot\frac{d\mathbf{e}(t)}{dt}$ (53)
Inserting Eq. (51) we have that
[…]
@@
where $\underline{\alpha}_{\Im}(\omega)$ is the the matrix defined in Eq. (7) of the main text. If we require that the energy dissipated during the perturbation is always positive irrespective of the form of the perturbation, then for $\omega>0$ the matrix $\underline{\alpha}_{\Im}(\omega)$ must be positive definite.
@@ Appendix B Symmetry properties
In this appendix we provide additional information about the consequences of the mirror-symmetry on the frictional quantum system discussed in Sec. III.
@@
If we consider the symmetry described by $\underline{M}_{n}$, Eq. (25) reveals that for the setup under consideration the above expression must coincide with
@@
Here, we have used that diagonal matrices and $\underline{\Lambda}_{n}$ are invariant with respect to the transformation while for the other matrices we have $\underline{M}_{n}\bm{\underline{\Lambda}}_{\perp}\underline{M}^{-1}_{n}=-\bm{\underline{\Lambda}}_{\perp}$ and $\underline{M}_{n}\mathbf{\underline{L}}_{\perp}\underline{M}^{-1}_{n}=-\mathbf{\underline{L}}_{\perp}$. When comparing the coefficients of Eqs. (B) and (B), we obtain $\mathbf{c}_{\perp}(-q,\mathbf{R}_{a},\omega)=-\mathbf{c}_{\perp}(q,\mathbf{R}_{a},\omega)$ (odd function in $q$) and $s_{n}(q,\mathbf{R}_{a},\omega)=s_{n}(-q,\mathbf{R}_{a},\omega)$ (even function in $q$). This is, however, in contradiction with the the constraints given by Eq. (24a) and, consequently, these expressions must vanish. This leads to the decomposition in Eq. (26). We can follow a similar procedure if the system features an additional symmetry $\underline{M}_{t}$, where $\mathbf{t}\cdot\mathbf{n}=0$. If the symmetry plane contains both, $\mathbf{q}=q\mathbf{n}$ and $\mathbf{R}_{a}$, thus making them invariant with respect to $\underline{M}_{t}$, we must have that $\underline{G}_{\Im}$ in Eq. (26) is invariant with respect to this transformation. However, if we call $\mathbf{t^{\prime}}=\mathbf{n}\times\mathbf{t}$, we obtain that
@@
$\underline{S}(\omega,v)$ $=\underline{\Delta}(\omega,v)+\bm{\chi}(\omega,v)\cdot\mathbf{\underline{\Lambda}}+\bm{\Omega}(\omega,v)\cdot\mathbf{\underline{L}}$
$=\underline{\Delta}(\omega,v)$
$\hskip 28.45274pt+\chi_{n}(\omega,v)\underline{\Lambda}_{n}+\bm{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}$
$\hskip 28.45274pt+\Omega_{n}(\omega,v)\underline{L}_{n}+\bm{\Omega}_{\perp}(\omega,v)\cdot\mathbf{\underline{L}}_{\perp}~{}.$ (60)
Due to the symmetry of the system and its dissipative nature, we can anticipate that $\Omega_{n}=0$. Arguments analogous to those made for $\underline{G}_{\Im}$ reveal that $\underline{\Delta}$ and $\chi_{n}$ are even functions of the velocity $v$, while $\bm{\chi}_{\perp}$ and $\bm{\Omega}_{\perp}$ must be a odd function of $v$. Therefore, we can write
$\underline{S}^{\sf T}(\omega,-v)$ $=\underline{\Delta}^{\sf T}(\omega,-v)+\bm{\chi}(\omega,-v)\cdot\mathbf{\underline{\Lambda}}^{\sf T}$
$\hskip 59.75095pt+\bm{\Omega}_{\perp}(\omega,-v)\cdot\mathbf{\underline{L}}_{\perp}^{\sf T}$
$=\underline{\Delta}(\omega,v)+\chi_{n}(\omega,v)\underline{\Lambda}_{n}-\bm{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}$
$\hskip 59.75095pt+\bm{\Omega}_{\perp}(\omega,v)\cdot\mathbf{\underline{L}}_{\perp}$
$=\underline{S}(\omega,v)-2\bm{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}~{},$ (61)
which is equivalent to Eq. (20). Again, when an additional mirror-symmetry $\underline{M}_{t}$ occurs and $\mathbf{v}=v\mathbf{n}$ and $\mathbf{R}_{a}$ are contained in (or parallel to) the symmetry plane, then $\underline{S}(\omega,v)$ must be invariant with respect to the transformation. With arguments analogous to the above-discussed case of $\underline{G}_{\Im}$, we can then conclude that $\bm{\chi}=\chi_{t}\mathbf{t}$ and $\bm{\Omega}_{\perp}=\Omega_{t}\mathbf{t}$.

### v2

@@ Appendix A Positive definite matrices
In this appendix, we elaborate on the reasons for why some of the matrices encountered in the main text are positive semidefinite or positive definite. We work in the Heisenberg picture in which the operator are time-dependent and the states are not. Let us consider first the power-spectrum tensor. For a generic complex vector $\mathbf{u}$ and any real-valued function $f(t)$ we can define the operator
$\hat{b}_{\mathbf{u}}(t)=\int_{-\infty}^{\infty}dt_{1}f(t-t_{1})\mathbf{u}\cdot\hat{\mathbf{d}}(t_{1})~.$ (43)
@@
$\langle\hat{b}^{{\dagger}}_{\mathbf{u}}(t)\hat{b}_{\mathbf{u}}(t)\rangle=\int_{-\infty}^{\infty}|f(\omega)|^{2}\mathbf{u}^{*}\cdot\underline{S}(\omega)\cdot\mathbf{u}\;d\omega\geq 0~.$ (44)
Then, since $|f(\omega)|^{2}\geq 0$, we must have that
@@
For discussing the properties of $\underline{\alpha}_{\Im}(\omega)$ we consider the perturbed Hamiltonian $\hat{\tilde{H}}_{t}=\hat{H}_{t}-\hat{\mathbf{d}}\cdot\mathbf{e}(t)$, where $\mathbf{e}(t)$ denotes a small ($\left|\mathbf{e}(t)\right|\ll 1$) real three-dimensional vector which is different from zero for a finite interval of time $\mathbf{e}(t\to\pm\infty)\to 0$. The perturbation introduces an additional intrinsic time-dependence with respect to that possibly characterizing the unperturbed Hamiltonian. In linear order of perturbation theory, the expectation value of the associated time evolution in the Heisenberg picture of the dipole operator as prescribed by $\hat{\tilde{H}}_{t}$ (see also below), i.e. $\hat{\tilde{\mathbf{d}}}(t)$, can be written as
@@
$\langle\hat{\tilde{\mathbf{d}}}(t)\rangle\equiv\mathrm{tr}[\hat{\tilde{\mathbf{d}}}(t)\hat{\rho}$] $\stackrel{{\scriptstyle\left|\mathbf{e}(t)\right|\ll 1}}{{\approx}}\int_{-\infty}^{\infty}\mathrm{d}t_{1}\;\underline{\alpha}(t-t_{1})\cdot\mathbf{e}(t_{1})~,$ (46)
@@
where we used the assumption that $\langle\hat{\mathbf{d}}(t)\rangle=0$ (see Sec. 2) and $\underline{\alpha}(\tau)$ is the linear response tensor with elements defined in Eq. (6). Mathematically, we may regard the tensor as the functional derivative with respect to the external perturbation of the exact time evolution, i.e. $\delta\langle\hat{\tilde{d}}_{i}(t)\rangle/\delta e_{j}(t_{1})=\alpha_{ij}(t-t_{1})$.
@@
In general, the time evolution according to $\hat{\tilde{H}}_{t}$ of an operator $\hat{A}_{t}$ which intrinsically depends on time is given by $\hat{\tilde{A}}(t)=\hat{\tilde{U}}^{{\dagger}}_{t}\hat{A}_{t}\hat{\tilde{U}}_{t}$,. The operator $\hat{\tilde{U}}_{t}$ is the solution of $\partial_{t}\hat{\tilde{U}}_{t}=(\mathrm{i}/\hbar)\hat{\tilde{H}}_{t}\hat{\tilde{U}}_{t}$ with the initial condition $\hat{\tilde{U}}_{t=0}=\underline{\mathds{1}}$. Accordingly, the total time derivative of $\hat{\tilde{A}}_{t}(t)$ is given by
$\frac{d}{dt}\hat{\tilde{A}}(t)=\frac{\mathrm{i}}{\hbar}\hat{\tilde{U}}^{{\dagger}}_{t}[\hat{\tilde{H}}_{t},\hat{A}_{t}]\hat{\tilde{U}}_{t}+\hat{\tilde{U}}^{{\dagger}}_{t}(\partial_{t}\hat{A}_{t})\hat{\tilde{U}}_{t}~.$ (47)
@@
We notice that for $\hat{A}_{t}\equiv\hat{\tilde{H}}_{t}$ the first term on the r.h.s. of the previous equation is identically zero. As a consequence, during the perturbation, the total change of energy is given by
$\int_{-\infty}^{\infty}dt\;\frac{d}{dt}\langle\hat{\tilde{H}}_{t}(t)\rangle$ $=\int_{-\infty}^{\infty}dt\;\langle\frac{d}{dt}\hat{\tilde{H}}_{t}(t)\rangle=\int_{-\infty}^{\infty}dt\;\langle U^{{\dagger}}_{t}(\partial_{t}\hat{\tilde{H}}_{t})U_{t}\rangle$
@@
$=\int_{-\infty}^{\infty}dt\;\langle U^{{\dagger}}_{t}(\partial_{t}\hat{H}_{t})U_{t}\rangle-\int_{-\infty}^{\infty}\mathrm{d}t\;\langle\hat{\tilde{\mathbf{d}}}(t)\rangle\cdot\frac{d\mathbf{e}(t)}{dt}~.$ (48)
While the first term in the last line of the previous expression gives the change in energy due to the intrinsic time dependence of the unperturbed Hamiltonian $\hat{H}_{t}$, the second term describes the modification of the energy due to the perturbation alone. Inserting Eq. (46) we have that
$\int_{-\infty}^{\infty}\mathrm{d}t\;\langle\hat{\tilde{\mathbf{d}}}(t)\rangle\cdot\frac{d\mathbf{e}(t)}{dt}=\int_{0}^{\infty}\frac{\mathrm{d}\omega}{\pi}\;\omega\;\mathbf{e}^{*}(\omega)\cdot\underline{\alpha}_{\Im}(\omega)\cdot\mathbf{e}(\omega)~,$ (49)
@@
where $\underline{\alpha}_{\Im}(\omega)$ is the the matrix defined in Eq. (7) of the main text. If we require that the change of energy induced by the perturbation is always positive irrespective of the form of the perturbation, then for $\omega>0$ the matrix $\underline{\alpha}_{\Im}(\omega)$ must be positive definite. Physically, this is equivalent to the request that, despite its intrinsic and possibly externally driven dynamics, the system remains stable and is able to dissipate the energy arising from the perturbation. Moreover, since [53]
$\underline{\alpha}(\mathrm{i}\xi)=\frac{1}{\mathrm{i}\pi}\int_{-\infty}^{\infty}d\omega\frac{\omega}{\omega^{2}+\xi^{2}}\underline{\alpha}(\omega)$ (50)
we can write that
$\frac{\underline{\alpha}(\mathrm{i}\xi)+\underline{\alpha}^{\sf T}(\mathrm{i}\xi)}{2}=\frac{1}{\pi}\int_{-\infty}^{\infty}d\omega\frac{\omega}{\omega^{2}+\xi^{2}}\underline{\alpha}_{\Im}(\omega)~.$ (51)
Given that the real part of a positive definite matrix is also positive definite and that the real and imaginary part of $\underline{\alpha}_{\Im}(\omega)$ are, respectively, odd and even in $\omega$, we obtain the expression in Eq. (8) of the main text. It indicates that, if $\underline{\alpha}_{\Im}(\omega)$ is positive definite, the symmetric part of $\underline{\alpha}(\mathrm{i}\xi)$ for $\xi\geq 0$ is also positive definite.
@@ Appendix B Symmetry properties
In this appendix, we provide additional information about the consequences of the mirror-symmetry on the frictional quantum system discussed in Sec. 3.
@@
If we consider the symmetry described by $\underline{M}_{n}=\underline{\mathds{1}}-2\mathbf{n}\mathbf{n}^{\sf T}$, Eq. (26) reveals that for the setup under consideration the above expression must coincide with
@@
Here, we have used that diagonal matrices and $\underline{\Lambda}_{n}$ are invariant with respect to the transformation while for the other matrices we have $\underline{M}_{n}\boldsymbol{\underline{\Lambda}}_{\perp}\underline{M}^{-1}_{n}=-\boldsymbol{\underline{\Lambda}}_{\perp}$ and $\underline{M}_{n}\mathbf{\underline{L}}_{\perp}\underline{M}^{-1}_{n}=-\mathbf{\underline{L}}_{\perp}$. When comparing the coefficients of Eqs. (B) and (B), we obtain $\mathbf{c}_{\perp}(-q,\mathbf{R}_{a},\omega)=-\mathbf{c}_{\perp}(q,\mathbf{R}_{a},\omega)$ (odd function in $q$) and $s_{n}(q,\mathbf{R}_{a},\omega)=s_{n}(-q,\mathbf{R}_{a},\omega)$ (even function in $q$). This is, however, in contradiction with the the constraints given by Eq. (25a) and, consequently, these expressions must vanish. This leads to the decomposition in Eq. (27). We can follow a similar procedure if the system features an additional symmetry $\underline{M}_{t}$, where $\mathbf{t}\cdot\mathbf{n}=0$. If the symmetry plane contains both, $\mathbf{q}=q\mathbf{n}$ and $\mathbf{R}_{a}$, thus making them invariant with respect to $\underline{M}_{t}$, we must have that $\underline{G}_{\Im}$ in Eq. (27) is invariant with respect to this transformation. However, if we call $\mathbf{t^{\prime}}=\mathbf{n}\times\mathbf{t}$, we obtain that
@@
$\underline{S}(\omega,v)$ $=\underline{\Delta}(\omega,v)+\boldsymbol{\chi}(\omega,v)\cdot\mathbf{\underline{\Lambda}}+\boldsymbol{\Omega}(\omega,v)\cdot\mathbf{\underline{L}}$
$=\underline{\Delta}(\omega,v)+\chi_{n}(\omega,v)\underline{\Lambda}_{n}+\boldsymbol{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}+\Omega_{n}(\omega,v)\underline{L}_{n}+\boldsymbol{\Omega}_{\perp}(\omega,v)\cdot\mathbf{\underline{L}}_{\perp}~.$ (57)
Due to the symmetry of the system and its dissipative nature, we can anticipate that $\Omega_{n}=0$. Arguments analogous to those made for $\underline{G}_{\Im}$ reveal that $\underline{\Delta}$ and $\chi_{n}$ are even functions of the velocity $v$, while $\boldsymbol{\chi}_{\perp}$ and $\boldsymbol{\Omega}_{\perp}$ must be a odd function of $v$. Therefore, we can write
$\underline{S}^{\sf T}(\omega,-v)$ $=\underline{\Delta}^{\sf T}(\omega,-v)+\boldsymbol{\chi}(\omega,-v)\cdot\mathbf{\underline{\Lambda}}^{\sf T}+\boldsymbol{\Omega}_{\perp}(\omega,-v)\cdot\mathbf{\underline{L}}_{\perp}^{\sf T}$
$=\underline{\Delta}(\omega,v)+\chi_{n}(\omega,v)\underline{\Lambda}_{n}-\boldsymbol{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}+\boldsymbol{\Omega}_{\perp}(\omega,v)\cdot\mathbf{\underline{L}}_{\perp}$
$=\underline{S}(\omega,v)-2\boldsymbol{\chi}_{\perp}(\omega,v)\cdot\mathbf{\underline{\Lambda}}_{\perp}~,$ (58)
which is equivalent to Eq. (21). Again, when an additional mirror-symmetry $\underline{M}_{t}$ occurs and $\mathbf{v}=v\mathbf{n}$ and $\mathbf{R}_{a}$ are contained in (or parallel to) the symmetry plane, then $\underline{S}(\omega,v)$ must be invariant with respect to the transformation. With arguments analogous to the above-discussed case of $\underline{G}_{\Im}$, we can then conclude that $\boldsymbol{\chi}=\chi_{t}\mathbf{t}$ and $\boldsymbol{\Omega}_{\perp}=\Omega_{t}\mathbf{t}$.
[…]